how do you solve r+4/3 = r/5 ? and problems similar to that one?
Answer by kiru_khandelwal(79)

r + 4/3 = r/5
=> (3r + 4)/3 = r/5
=> 5(3r + 4) = 3*r
=> 15r + 20 = 3r
=> 15r - 3r = -20
=> 12r = -20
=> r = -20/12
=> r = -5/3